Test the hypothesis. Show the step-by-step process in testing hypothesis.



1. A researcher claims that the average salary of a private school teacher is greater than P35,000 with a standard deviation of P7,000. A sample of 35 teachers has a mean salary of P37,000. Test the claim of the researcher. At 0.05 level of significance, test the claim of the researcher.



2. A researcher reports that the average salary of a college dean is more than P65,000. A sample of 35 college deans has a mean salary of P67,000. At 0.01 level of significance, test the claim that the college deans earn more than P65,000 a month. The standard deviation of the population is P5,250.

1
Expert's answer
2022-04-27T12:51:16-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu\le35000"

"H_1:\mu>35000"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05," "df=n-1=35-1=34" degrees of freedom, and the critical value for a right-tailed test is "t_c = 1.690924."

The rejection region for this right-tailed test is "R = \{t: t> 1.690924\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{37000-35000}{7000/\sqrt{35}}\approx1.6903"

Since it is observed that "t =1.6903 <1.690924=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=34" degrees of freedom, "t=1.6903" is "p=0.05006," and since "p=0.05006>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 35000, at the "\alpha = 0.05" significance level.


2. The following null and alternative hypotheses need to be tested:

"H_0:\mu\le65000"

"H_1:\mu>65000"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z: z> 2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{67000-65000}{5250/\sqrt{35}}\approx2.2537"

Since it is observed that "z =2.2537<2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "z=2.2537" is "p=P(Z>2.2537)=0.012108," and since "p=0.012108>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 65000, at the "\alpha = 0.01" significance level.